OpenStudy (anonymous):

pq^2 pq^5 what is the greatest common factor of the monomials

OpenStudy (jango_in_dtown):

Let a=pq^2 and b=pq^5 then a=p^1q^2 and b=p^1q^5. gcd(a,b)=p^min{1,1} q^min{2,5}=p^1q^2=pq^2

OpenStudy (jango_in_dtown):

This is the generalization.@perryshann